# Stacked conformal prediction

### v2

@@ 2 Symmetric stack
Let $\hat{\psi}_{n}^{(y)}$ be a meta-learner trained from $\{(Z_{1},Y_{1}),\dots,(Z_{n},Y_{n}),(Z_{n+1},y)\}$, for $y\in\mathbb{R}$, and suppose the order of the $n+1$ pairs in this sample is irrelevant for the construction of $\hat{\psi}_{n}^{(y)}$. For a conformity function $\rho:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, define the conformity scores $R_{i}^{(y)}=\rho(Y_{i},\hat{\psi}^{(y)}_{n}(Z_{i}))$, for $i=1,\dots,n+1$, and let $R^{(y)}_{(1)}\leq R^{(y)}_{(2)}\leq\dots\leq R^{(y)}_{(n)}$ denote the ordered conformity scores among $\{R^{(y)}_{1},R^{(y)}_{2}\dots,R^{(y)}_{n}\}$. Choosing a nominal miscoverage level $10<\alpha<$ 1, we have that
@@ 3 Feasible stack
Given $\epsilon>0$, if there is a $\delta=\delta(\epsilon)>0$, such that
@@ 11             $y_{0}\leftarrow(\texttt{lower}+\texttt{inf})/2$
12             Get scores $r_{0}$ and $\hat{r}$ inputing $(z_{0},y_{0},Z,\alpha,A,\hat{\beta},B)$ to Algorithm 2
@@ 24             $y_{0}\leftarrow(\texttt{sup}+\texttt{upper})/2$
25             Get scores $r_{0}$ and $\hat{r}$ inputing $(z_{0},y_{0},Z,\alpha,A,\hat{\beta},B)$ to Algorithm 2
@@ 33      $c_{i,2}\leftarrow\texttt{sup}$
Algorithms 1 and 2 formally describe the efficient conformalization of a multiple linear regression meta-learner based on this classical formula (see line 1 of Algorithm 1). It’s worth mentioning that the residuals determination in Algorithm 2 do not involve the computation of additional inverses. The use of these residuals to construct the conformity scores denominators is crucial for obtaining prediction intervals with more adaptive width.
@@ 5 Concluding remarks and open source implementations
We have presented the stacked conformal prediction procedure in a regression setting, but its general structure readily extends to other tasks – classification being a prime example. The key requirement in any such extension is the choice of a meta-learner that admits full conformalization at manageable computational cost. In classification problems, a natural meta-learner candidate would be the $k$-nearest neighbors classifier, equipped with a distance metric tailored to aggregate the class-probability outputs produced by the stack base-learners. R (R Core Team, 2024) and Python (Python Core Team, 2023) codes, including C++ (Stroustrup, 2013) extensions, for full reproduction of the examples in the paper are available at:
@@
Let $U_{1},U_{2},\dots,U_{n}$ be exchangeable random variables, and denote their order statistics by ${U_{(1)}\leq U_{(2)}\leq\dots\leq U_{(n)}}$. It follows that $P(U_{i}\leq U_{(k)})\geq k/n$, for each ${i,k=1,\dots,n}$.
@@
Considering that we may have ties, it follows from the definition of $U_{(k)}$ that at least $k$ of the $U_{i}$’s are less than or equal to $U_{(k)}$. Hence, ${\sum_{i=1}^{n}I_{\{U_{i}\leq U_{(k)}\}}\geq k}$, almost surely. Taking expectations and noting that, by exchangeability, the probability $P(U_{i}\leq U_{(k)})$ is the same for every ${i=1,\dots,n}$, the result follows.
@@
for $k=1,\dots,n+1$. Choosing $k=\lceil(1-\alpha)(n+1)\rceil\geq(1-\alpha)(n+1)$, the result follows, since the definition of the random prediction set $C_{n+1}^{(\alpha)}$ entails that $Y_{n+1}\in C_{n+1}^{(\alpha)}$ if and only if $R^{(Y_{n+1})}_{n+1}\leq R^{(Y_{n+1})}_{(\lceil(1-\alpha)(n+1)\rceil)}$.

### v3

@@ 2 Symmetric stack
Let $\hat{\psi}_{n}^{(y)}$ be a meta-learner trained from $\{(Z_{1},Y_{1}),\dots,(Z_{n},Y_{n}),(Z_{n+1},y)\}$, for $y\in\mathbb{R}$, and suppose the order of the $n+1$ pairs in this sample is irrelevant for the construction of $\hat{\psi}_{n}^{(y)}$. For a conformity function $\rho:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$, define the conformity scores $R_{i}^{(y)}=\rho(Y_{i},\hat{\psi}^{(y)}_{n}(Z_{i}))$, for $i=1,\dots,n+1$, and let $R^{(y)}_{(1)}\leq R^{(y)}_{(2)}\leq\dots\leq R^{(y)}_{(n)}$ denote the ordered conformity scores among $\{R^{(y)}_{1},R^{(y)}_{2}\dots,R^{(y)}_{n}\}$. Choosing a nominal miscoverage level $10<\alpha<$ 1 such that $\lceil(1-\alpha)(n+1)\rceil\leq n$, we have that
@@ 3 Feasible stack
Given $\epsilon>0$, if there is a $\delta=\delta(\epsilon)>0$ such that
@@ 11             $y_{0}\leftarrow(\texttt{lower}+\texttt{inf})/2$
12             Get scores $r_{0}$ and $\hat{r}$ inputing $(z_{0},y_{0},Z,y,\alpha,A,\hat{\beta},B)$ to Algorithm 2
@@ 24             $y_{0}\leftarrow(\texttt{sup}+\texttt{upper})/2$
25             Get scores $r_{0}$ and $\hat{r}$ inputing $(z_{0},y_{0},Z,y,\alpha,A,\hat{\beta},B)$ to Algorithm 2
@@ 33      $c_{i,2}\leftarrow\texttt{sup}$
Algorithms 1 and 2 formally describe the efficient conformalization of a multiple linear regression meta-learner based on this classical formula (see line 1 of Algorithm 1). It is worth mentioning that the residuals determination in Algorithm 2 do not involve the computation of additional inverses. The use of these residuals to construct the conformity scores denominators is crucial for obtaining prediction intervals with more adaptive width.
@@ 5 Concluding remarks and open source implementations
We have presented the stacked conformal prediction procedure in a regression setting, but its general structure readily extends to other tasks — classification being a prime example. The key requirement in any such extension is the choice of a meta-learner that admits full conformalization at manageable computational cost. In classification problems, a natural meta-learner candidate would be the $k$-nearest neighbors classifier, equipped with a distance metric tailored to aggregate the class-probability outputs produced by the stack base-learners. R (R Core Team, 2024) and Python (Python Core Team, 2023) code, including C++ (Stroustrup, 2013) extensions, for full reproduction of the examples in the paper are available at:
@@
Let $U_{1},U_{2},\dots,U_{n},U_{n+1}$ be exchangeable random variables, and denote the order statistics of the subset $\{U_{1},U_{2},\dots,U_{n}\}$ by $V_{(1)}\leq V_{(2)}\leq\dots\leq V_{(n)}$. For each ${k=1,\dots,n}$, it follows that $P(U_{n+1}\leq V_{(k)})\geq k/(n+1)$.
@@
Denote the order statistics of the full set $\{U_{1},U_{2},\dots,U_{n},U_{n+1}\}$ by
$U_{(1)}\leq U_{(2)}\leq\dots\leq U_{(n)}\leq U_{(n+1)}.$
Considering that we may have ties, it follows from the definition of $U_{(k)}$ that at least $k$ of the $U_{i}$’s are less than or equal to $U_{(k)}$, for $k=1,\dots,n+1$. Hence, ${\sum_{i=1}^{n+1}I_{\{U_{i}\leq U_{(k)}\}}\geq k}$, almost surely. Taking expectations and observing that, by exchangeability, the probability ${P(U_{i}\leq U_{(k)})}$ is the same, for ${i=1,\dots,n+1}$, we have that ${P(U_{n+1}\leq U_{(k)})\geq k/(n+1)}$. Furthermore, for $k=1,\dots,n$, notice that $U_{n+1}>V_{(k)}$ if and only if $U_{n+1}>U_{(k)}$, because $U_{n+1}$ cannot be strictly larger than itself. Hence, ${P(U_{n+1}\leq V_{(k)})=P(U_{n+1}\leq U_{(k)})}$, for $k=1,\dots,n$, and the result follows.
@@
for $k=1,\dots,n$. Choosing $k=\lceil(1-\alpha)(n+1)\rceil\geq(1-\alpha)(n+1)$, the result follows, since the definition of the random prediction set $C_{n+1}^{(\alpha)}$ entails that $Y_{n+1}\in C_{n+1}^{(\alpha)}$ if and only if $R^{(Y_{n+1})}_{n+1}\leq R^{(Y_{n+1})}_{(\lceil(1-\alpha)(n+1)\rceil)}$.